A number X is selected at random from the unit interval. Let the events A and B be:
A = “X differs from 1/2 by more than 1/4”
B = “1 - X is less than 1/2.”
Find the events A∩B, A^c∩B, A∪B
1
Expert's answer
2019-07-24T13:46:48-0400

Set A =


"A=\{|x-1/2|>1/4\}="



"\{x-1/2>1/4\ or\ x-1/2<-1/4\}="

"\{x>3/4\ or\ x<1/4\}="


"(0,\frac{1}{4})\cup (\frac{3}{4},1)"

Set B =


"B=\{1-x<\frac{1}{2}\}=\{x>\frac{1}{2}\}=(\frac{1}{2},1)"

Set A∩B =

"A\cap B=((0,\frac{1}{4})\cup (\frac{3}{4},1))\cap(\frac{1}{2},1)="

"((0,\frac{1}{4})\cap (\frac{1}{2},1))\cup((\frac{3}{4},1)\cap (\frac{1}{2},1))="

"(\frac{3}{4},1)"

Set Ac =


"A^c=((0,\frac{1}{4})\cup (\frac{3}{4},1))^c"

"[\frac{1}{4},\frac{3}{4}]"

Set Ac∩B =

"[\frac{1}{4},\frac{3}{4}]\cap (\frac{1}{2},1)=(\frac{1}{2},\frac{3}{4}]"

Set A∪B =


"((0,\frac{1}{4})\cup (\frac{3}{4},1))\cup (\frac{1}{2},1)="

"(0,\frac{1}{4})\cup((\frac{3}{4},1)\cup (\frac{1}{2},1))="

"(0,\frac{1}{4})\cup(\frac{1}{2},1)"

P(A∩B) = 1-3/4=1/4, P(Ac∩B) = 3/4-1/2 = 1/4, p(A∪B) = (1/4 - 0) + (1 - 1/2) = 3/4.

Answer:A∩B = (3/4, 1) , Ac∩B = (1/2, 3/4], A∪B = (0, 1/4)∪(1/2, 1).